### v1

# Two-mode squeezing and SU(1,1) interferometry with trapped ions

Abstract
We experimentally implement circuits of one and two mode operations on two motional modes of a single trapped ion. This is achieved by implementing the required displacement, squeezing, two-mode squeezing, and beamsplitter operations using oscillating electric potentials applied to the trap electrodes. The resulting electric fields drive the modes resonantly or parametrically without the need for optical forces. As a demonstration, we implement SU(2) and SU(1,1) interferometers with phase sensitivities near the Cramér-Rao bound. We report a maximum sensitivity of a SU(2) interferometer within 0.67(5) dB of the standard quantum limit (SQL) as well as a single and two-mode SU(1,1) sensitivity of 5.9(2) dB and 4.5(2) dB below the SQL respectively.
preprint: APS/123-QED
Quantum-mechanical harmonic oscillators, described by infinite dimensional, non-commuting observables, are currently being utilized for quantum sensing [1, 2] and studies of continuous variable quantum computing (CVQC) [3]. The quantum-mechanical harmonic oscillator is realized in trapped-ions, nano-mechanical systems [4] and has an analog with microwave photons in superconducting circuits [5] and optical photons in photonic systems [6]. Recent work, using motional states of trapped-ions has shown the ability to detect small displacements [7, 8] and encode logical qubits [9, 10]. The use of two-mode squeezed states could provide a quantum advantage [11](larger quantum signal to noise ratio with the same number of resources) for sensing and significant work has been done to generate these states in other platforms [12, 13, 14]. Two-mode squeezed states have already proven useful for quantum sensing using SU(1,1) interferometers [15, 16, 17, 18] in photonics, but can be difficult to generate efficiently and suffer from photon loss.
In the following, we demonstrate different classes of interferometers using the motional modes of a single trapped atomic ion. With a demonstration of these interferometers, we show the ability to concatenate several operations on the oscillator states of two modes in a phase coherent manner. Performing these operations with digitally programmable frequency sources for relative phase control simplifies significant calibration challenges when developing more complicated circuits.
Trapped ions are well suited for such demonstrations due to the harmonic nature of the trap and the potential to achieve relatively long mechanical oscillator coherence times [19]. Reconstructing the motional states of these oscillators is achieved by coupling them to the ion’s internal states via motional sideband transitions and measuring the internal state populations [20]. In the limit of a small number of motional quanta these transitions can approximate a phonon counting type measurement. This enables a homodyne type detection scheme which is used to quantify the sensitivity we achieve with the different classes of interferometers demonstrated below.
Yurke et al. [21] proposed schemes to reach the Heisenberg limit [22] using different classes of interferometers. Here, we work within the framework of the Cramér-Rao bound [23, 24] which sets the ultimate limit for the resolution of a sensing scheme. This bound is described by a parameter variance, instead of the Heisenberg limit which is set by the number of resources used, i.e. photon or phonon number [24], making it more suitable to the limits of the experiments reported here. The required interactions for the interferometers (described below) are colloquially called displacing, beam splitting, single-mode squeezing, and two-mode squeezing. For trapped ions, these interactions can be implemented with laser beams [25] or in a laser-free method using additional potentials applied to the trap electrodes [26, 27]. Performance of these interferometers can, in practice, be limited only by the motional heating of the ion resulting from fluctuating patch potentials on the trap electrodes [28, 29]. Here, we demonstrate an SU(2) interferometer (Mach-Zender), which has previously been demonstrated in [25, 30, 31], (coherent input states are used here in contrast to these works), as well as single-mode and two-mode SU(1,1) [^1] The Lie-group labelling is used to refer to the groups that contain the operators that generate these states. interferometers. Each of the interferometers described here is sensitive to different phase shifts, which are single-mode phase $\phi_{a}$ (single-mode SU(1,1)), difference phase $\phi_{a}-\phi_{b}$ (SU(2)), or sum phase $\phi_{a}+\phi_{b}$ (two-mode SU(1,1)), where the $a,b$ labels refer to different modes. We show performance close to the Cramér-Rao bounds for these interferometers.
To demonstrate these different interferometers we use a single trapped ${}^{40}$${}^{+}$ ion as the mechanical oscillator. The ion is confined in a macroscopic, room-temperature, linear Paul trap with an ion-electrode spacing of $\simeq$ 0.75 mm. Experiments are performed using the two ‘radial’ modes of the ion (normal to the trap axis) with frequencies $\omega_{a}\approx 2\pi\times 1.80$ MHz and $\omega_{b}\approx 2\pi\times 1.83$ MHz, and oscillator energy eigenstates denoted by |$\ket{n_{a},n_{b}}$. The mode splitting can be adjusted by changing a DC potential on the RF electrodes [32] which also rotates the mode angle and can be used to vary the achievable coupling rate of the interactions described below. We use qubit states ||$\ket{\uparrow}\equiv\ket{m_{L}=+5/2}$ and ||$\ket{\downarrow}\equiv\ket{m_{L}=+3/2}$ within the metastable $D_{5/2}$ manifold which has a lifetime of $\simeq$ 1.1 seconds [33], where $m_{L}$ is the total angular momentum projection along the direction of the quantization magnetic field of 0.498 G. The qubit transition frequency is $\omega_{0}\approx 2\pi\times 2.63$ MHz. This is a magnetic field sensitive qubit with a coherence time of $\simeq$ 1  ms, while stable trapping potentials provide motional coherence times over 20 ms. The qubit is manipulated using an orthogonal pair of laser beams red detuned from the transition frequency between $P_{3/2}$ and $D_{5/2}$ manifolds by 44 THz to induce coherent stimulated-Raman transitions between the two qubit states. In each experiment the qubit is first prepared in the ||$\ket{\uparrow}\ket{0,0}$ state with high fidelity, using optical pumping, electromagnetically-induced-transparency (EIT) cooling [34] implemented with a pair of 397 nm beams on the $S_{1/2}\rightarrow P_{1/2}$ transition and resolved-sideband laser cooling using the qubit described above. Mode temperatures correspond to $\bar{n}_{a,b}<0.1$. Qubit readout is performed by “deshelving” population in the $\ket{\downarrow}$ state followed by applying a laser beam resonant with the $S_{1/2}\leftrightarrow P_{1/2}$ cycling transition and detecting state-dependent ion fluorescence [35]. Motional state analysis is performed by setting the Raman beams difference frequency to $\omega_{0}\pm\omega_{m}$, where $m$ denotes the desired mode for readout. Applying these sideband interactions maps the motion onto the qubit states and produces Rabi oscillations with multiple frequency components having amplitudes that depend on the Fock state populations [20]. This interaction is generated by applying the same pair of Raman beams that drive the qubit, but detuned above (red sideband) or below (blue sideband), the qubit frequency. In the Lamb-Dicke limit [20], the red sideband interaction drives transitions between ||$\ket{\uparrow}\ket{n}$ and ||$\ket{\downarrow}\ket{n-1}$ with Rabi frequencies proportional to $\sqrt{n}$ and can be used to verify preparation or return to the motional ground state. The blue sideband interaction drives transitions between ||$\ket{\uparrow}\ket{n}$ and ||$\ket{\downarrow}\ket{n+1}$ with Rabi frequencies proportional to $\sqrt{n+1}$ and is used to fully characterize the generated state populations [20].
Figure 1: (a) View of trap along the trap axis showing the electrical connections:pink—RF quadrupole electrodes, trap and parametric drive, purple—displacement drive and DC electrode, yellow— other DC electrodes and relevant control electrodes. The green box represents the circuitry necessary to combine the drives [32]. Mode $a$ and mode $b$ label the two different ‘radial’ mode vectors of a single ion with the angles from the horizontal shown. (b) A 3D rendering of the trap showing direction of the 976 nm laser used for motional state preparation and state characterization, and cyan ball representing the ${}^{40}$${}^{+}$ ion
Generation of the motional states is accomplished by applying time-varying potentials to the trap electrodes (see Fig. 1) producing a unitary operation. The displacement operation is produced by applying an oscillating potential, at the mode frequency, to a single trap electrode (purple electrode Fig. 1). Single-mode squeezing, two-mode squeezing and the beam splitter are accomplished by applying an oscillating potential at twice the mode frequency $(2\omega_{a})$, at the sum of the mode frequencies $(\omega_{a}+\omega_{b})$ or at the difference frequency $(\omega_{a}-\omega_{b})$ respectively (pink electrodes Fig. 1). These modulated potentials yields a generalized squeezing interaction Hamiltonian,
$\hat{H}=i\hbar\dfrac{g}{2}\big{[}\hat{a}_{i}\hat{a}_{j}\exp(i\delta t)\exp(-i\theta)-\hat{a}^{\dagger}_{i}\hat{a}^{\dagger}_{j}\exp(-i\delta t)\exp(i\theta)\big{]}$ (1)
where $g$ is the parametric coupling strength, which has a different value for each operation, $\delta$ is the detuning from the interaction resonance (either single mode squeezing (mode $a$ or $b$) or two-mode squeezing), and $\theta$ is the phase of the parametric modulation. With the proper choice of detuning this implements the unitary squeezing operator with squeezing parameter magnitude given by $r_{SMS}=gt$ and the two mode squeezing operator with $r_{TMS}=2gt$. Here, we used a single-mode displacement coupling of $g=2\pi\times 1.37$ kHz which gives a displacement parameter $\alpha=gt$. The maximum single mode squeezing coupling is $g=2\pi\times 3.99$ kHz and a maximum two-mode squeezing coupling of $g=2\pi\times 1.15$ kHz. Maximum coupling refers to a coupling rate calculated from fits to the states with the maximum voltage we used and the relevant pulse time for each operation. The beam splitter coupling is $g=2\pi\times 0.66$ kHz calculated using calibration data shown in [32].
In order to verify the proper performance of the beam splitter operation, we use a calibration experiment described in [32]. Verification of the two-mode squeezed state is more complicated given we are only able to readout a single mode in a single shot of the experiment. Using the blue sideband readout method, described above, is equivalent to tracing over a single mode and a thermal state is measured [36]. This is analogous to measuring a single qubit from a Bell state which produces a random qubit state. Additionally, we can use a beam splitter operation after the two-mode squeezer to disentangle the modes and produce single-mode squeezed states (Fig. 2). Performing both of these measurements verifies we have produced the target two-mode squeezed state and enables the use of these states for interferometric measurements.
Figure 2: (a) General quantum circuit diagram for characterizing the two-mode squeezed state with blue sideband readout performed at one of two instances, to verify we are producing the expected output (b) Measurement after the two-mode squeezed state and before beam splitter produces a thermal state. (c) applying the beam splitter operation after the two-mode squeezer and then characterizing the state produces two single-mode squeezed states with squeezing parameter $r$ equal to that of the two-mode squeezed state.
Implementing time reversal protocols is a method for generating these types of interferometers and is a powerful tool for sensing applications [37]. These protocols also enable us to verify that we are coherently generating the desired states by time-reversing back to the initial state. The SU(2) interferometer (Fig. 3c) is implemented by first displacing mode $a$ in our experiment, which generates the desired initial coherent state. The beam splitter is then used to generate a superposition of coherent and vacuum states. The phase of the final beam-splitter is controlled by using a variable delay, such that the accumulated phase is $(\omega_{a}-\omega_{b})t_{delay}$. The final beam splitter will completely transfer the phonons to mode $b$ with the correct choice of relative phase, and the red sideband readout will verify we have returned to the ground state (Fig. 3c). We use a maximum displacement of $\alpha\simeq\lvert 6\rvert$ to maintain a similar maximum $\bar{n}$ in the single and two-mode squeezed states.
The single-mode (Fig. 3a) and two-mode (Fig. 3b) SU(1,1) interferometers are implemented by using the parametric drive tuned to either twice the frequency of mode $a$ or the sum frequency of mode $a$ and $b$ respectively. To avoid increased decoherence, from long pulse durations, these pulses are run at a fixed duration with the drive voltage being varied to control the size of the state. A second squeezing pulse is then applied (Fig. 3b) with a varied phase controlled by the driving electronics. We are limited in the amount of two-mode squeezing we can generate due to our increased sensitivity to common-mode phase fluctuations that this interferometer is sensitive to.
Figure 3: (a) Single-mode SU(1,1) interferometer output for different amounts of squeezing is shown by varying the relative phase of the second squeezing pulse shown in the circuit diagram (inset). (b) repeat of (a) with two-mode squeezing and two-mode SU(1,1) shown in the interferometer circuit inset.(c) SU(2) interferometer output shown for various sizes of initial displacement where relative phase is controlled with a variable delay shown in the SU(2) circuit inset.
In order to determine the sensitivities to small phase shifts that are achievable with these different experiments, we calculate the quantum Fisher information for a pure state [23]
$\mathcal{F}(\phi)=4(\Delta\hat{\mathcal{G}})^{2}$ (2)
where $(\Delta\hat{\mathcal{G}})^{2}$ is the variance in the initial state of the probe (coherent state, squeezed state, two-mode squeezed state) and $\hat{\mathcal{G}}=i[d\hat{U}^{\dagger}/d\phi]\hat{U}$. For phase sensing $\hat{U}=e^{i\phi a^{\dagger}a}$ and therefore $\hat{\mathcal{G}}=a^{\dagger}a$ and the Fisher information is proportional to variance in the phonon number of the probe states. The phase sensitivity is limited by the Cramér-Rao bound [23] given by
$\Delta\phi\geq\dfrac{1}{\sqrt{\mathcal{F(\phi)}}}.$ (3)
This means for the SU(2) interferometer we are limited at
$\Delta\phi=\dfrac{1}{\alpha}=\dfrac{1}{\sqrt{\expectationvalue{N}_{d}}}$ (4)
where $\expectationvalue{N}_{d}$ is the expected mean phonon number for the displacement generated and this limit corresponds to the expected standard quantum limit. The limit for the single-mode SU(1,1) interferometer is given by [38]
$\Delta\phi=\dfrac{1}{\sqrt{8\expectationvalue{N}_{SMS}(\expectationvalue{N}_{SMS}+1)}}$ (5)
where $\expectationvalue{N}_{SMS}$ is the expected mean phonon number for the input squeezed state. The two-mode SU(1,1) limit is given by [23]
$\Delta\phi=\dfrac{1}{\sqrt{\expectationvalue{N}_{TMS}(2+\expectationvalue{N}_{TMS})}}$ (6)
where $\expectationvalue{N}_{TMS}$ is the expected mean phonon number for the input two-mode squeeze state.
The sensitivities achieved are calculated from the experimental data using the Fisher information [23]
$F(\phi)=\sum_{j}\dfrac{1}{P_{j}(\phi)}\bigg{[}\dfrac{dP_{j}(\phi)}{d\phi}\bigg{]}^{2}$ (7)
where $P_{j}(\phi)$ is the probability of getting the experimental results $j$ if the value of the parameter is $\phi$. In our experiment with only two possible outcomes, spin-up or spin-down, this simplifies to
$F(\phi)=\dfrac{1}{P_{\downarrow}(1-P_{\downarrow})}\bigg{[}\dfrac{\partial P_{\downarrow}}{\partial\phi}\bigg{]}^{2}.$ (8)
[…]
Figure 4: Ideal experimental phase sensitivities ($\delta\phi$) for the SU(2) (dashed green line), single mode SU(1,1) (dashed orange line) and two-mode SU(1,1) dashed red line) are shown and can be compared to experimental results shown as dots with error bars and Cramér-Rao (CR) bounds in solid lines.
We have demonstrated a wide class of trapped-ion motional state interferometers using a ‘laser-free’ parametric drive achieving single and two-mode SU(1,1) interferometer sensitivities of 5.9(2) dB and 4.5(2) dB below the SQL, respectively and a SU(2) interferometer sensitivity within 0.67(5) dB of the SQL. These results illustrate the measurement of phase shifts near the Cramér-Rao bound for common-mode, relative, and single-mode phase fluctuations. Enhanced sensitivity for measuring these motional phase fluctuations could enable characterization and mitigation of motional frequency noise which can benefit experiments involving motional states. Beside providing a framework for implementing quantum-enhanced sensing protocols, our methods provide the flexibility to concatenate multiple motional operations into circuits for potential applications in CVQC.
During preparation of this manuscript we became aware of related work using reservoir engineering to prepare two-mode squeezed states of trapped ions [40].
Acknowledgements.
This research is supported by support from NSF through the Q-SEnSE Quantum Leap Challenge Institute, Award #2016244, and the US Army Research Office under award W911NF-20-1-0037. The data supporting the figures in this article are available upon reasonable request from D.T.C.A.